### v1

# CloudHeatMap: Heatmap-Based Monitoring for Large-Scale Cloud Systems

Abstract
Cloud computing is essential for modern enterprises, requiring robust tools to monitor and manage Large-Scale Cloud Systems (LCS). Traditional monitoring tools often miss critical insights due to the complexity and volume of LCS telemetry data. This paper presents CloudHeatMap, a novel heatmap-based visualization tool for near-real-time monitoring of LCS health. It offers intuitive visualizations of key metrics such as call volumes, response times, and HTTP response codes, enabling operators to quickly identify performance issues. A case study on the IBM Cloud Console demonstrates the tool’s effectiveness in enhancing operational monitoring and decision-making. A demonstration is available at https://www.youtube.com/watch?v=3u5K1qp51EA.
Index Terms:
Cloud computing, Heatmaps, Cloud monitoring
I Introduction
Cloud computing has become a critical infrastructure for businesses, providing scalability and cost reduction. However, as Cloud adoption grows, monitoring these systems becomes increasingly complex [1, 2, 3]. Large-Scale Cloud Systems (LCS) comprise numerous interconnected microservices distributed across data centers (see [4] for an example), generating vast amounts of telemetry data [5] essential for understanding system health and performance. The sheer volume and complexity of these data challenge traditional monitoring tools, which often fail to provide timely, actionable insights [1, 2, 3]. This lack can lead to undetected performance issues, inefficient resource allocation, and, ultimately, system failures that impact service delivery [1, 2, 3].
Traditional Cloud monitoring tools (e.g., DataDog, Dynatrace) offer limited root cause visibility and are often reactive, alerting after issues impact the system. In dynamic Cloud environments, delayed responses can cause significant downtime and resource inefficiency. According to Google’s Site Reliability Engineering (SRE) principles, the four Golden Rules of Monitoring — traffic, errors, latency, and saturation — are essential metrics for assessing LCS health [6]. However, most monitoring tools focus on the interface-level analysis of system components [7], detecting issues but not providing deeper insights or predicting problems. While some advanced monitoring systems offer visual dashboards [8], they often lack support for exploratory health analysis. Operators need to delve into data patterns and understand component behaviours under varying conditions [9]. In today’s dynamic Cloud environments, early and proactive detection are critical [10, 4, 11, 12, 13, 14]; reactive approaches are insufficient. Thus, advanced monitoring solutions are needed that provide real-time insights and enable exploratory analysis to help Cloud Operators (Ops) teams identify issues before they escalate [1, 2, 3].
The authors, drawing from their experience designing and using AIOps tools [1, 2, 3, 10, 4], can attest that while these tools are useful, they have limitations in root cause analysis and detecting persistent issues. Long-standing abnormalities may escape automated detection, requiring human operators to assess the broader system and make strategic decisions beyond local failures.
The need arises for a mechanism that allows for a complete, bird’s eye view of a complex system — one that human cognition can easily process. The proposed tool addresses these gaps by providing near-real-time, intuitive visualizations for deeper insights into system health, enabling operators to diagnose and prevent issues before they affect system reliability. We are guided by the following research questions (RQs):
RQ1: How can we design visualizations that effectively support LCS monitoring and provide actionable insights?
RQ2: How do components across data centers respond to varying workloads, and how can we visualize these responses?
RQ3: How can we detect and visualize component issues in near-real-time for proactive management?
RQ4: How does our visualization tool impact LCS maintenance and operations?
We present CloudHeatMap [^1] The tool is available via https://github.com/sohanasarah/CloudHeatMap and includes $\approx 24$ hours of the anonymized real data., a novel heatmap-based [^2] A grid heatmap maps data points as colours in two dimensions, where colour intensity represents data magnitude [15], helping identify patterns and relationships in large datasets, allowing for quicker decision-making. visualization tool leveraging telemetry data from microservices to visualize key performance metrics like call volumes, response times, and HTTP (as well as custom) response codes.
We drew inspiration from the classical annunciator panels [16], that were commonly used in power plants and aircraft control boards. These panels feature a grid of lights or buttons, each representing a system parameter. When a parameter changes, the light or button updates its color or state, enabling operators to efficiently monitor processes and respond to alerts.
As we empirically found, heatmaps are valuable because they condense large amounts of data from complex systems into a simple grid, where color highlights anomalies, allowing operators to quickly and intuitively detect unusual behaviour.
CloudHeatMap offers a near-real-time system health overview, enabling operators to quickly identify hot spots and potential issues. The tool helps Cloud Ops make informed decisions about resource allocation, configuration changes, and problem resolution, ultimately enhancing Cloud services’ reliability and efficiency. It also supports temporal analysis, helping operators track system behaviour over time.
To evaluate tools’s effectiveness, we conducted a case study on the IBM Cloud Console, an LCS with significant operational demands. Results show that CloudHeatMap  enhances Cloud Ops’ ability to monitor and maintain system health, providing insights previously difficult to obtain with traditional methods. This research contributes to Cloud computing by offering a scalable and practical solution for monitoring LCS.
II Methodology
This research methodology [^3] For further details, please refer to the M.Sc. thesis by the first author [17]. involves the design, development, and evaluation of CloudHeatMap  tailored for monitoring the health of an LCS. The process is divided into key phases: system analysis, data collection, data processing, and visualization design. The diagram in Fig. 1 shows the sequence of actions required from data collection to heatmap generation.
Note that CloudHeatMap focuses on visualization (Section II-D) and operates independently of data collection and processing (Sections II-B and II-C). These phases are included as part of a representative example to demonstrate a complete data pipeline for context. Consequently, CloudHeatMap can be used with any data source, as long as the data is provided in the expected format.
Figure 1: System diagram.
II-A System Analysis
The first step involved analyzing the LCS under study, the IBM Cloud Console, a critical component of IBM’s Cloud infrastructure [4], serving as the web front-end and orchestrator for the IBM Cloud platform [18]. Currently, the system consists of around 125 microservices deployed worldwide. Each microservice generates millions of log records daily, encompassing various operational metrics, including HTTP response codes, call volumes, and response times [4].
Understanding the architecture and operational characteristics of the IBM Cloud Console was crucial for identifying appropriate metrics and effective visualization techniques. The complex distributed nature of the system and the volume of telemetry data posed significant challenges, making it ideal for testing the proposed visualization tool.
II-B Data Collection
The data collection phase focused on capturing telemetry data from the IBM Cloud Console. In this context, telemetry data refers to the metrics, logs, and traces emitted by the IBM Cloud Console microservices. The traces contain detailed information about each request processed by the microservices, including timestamps, response codes, response times, and the identities of caller and callee microservices [^4] A user request triggers calls between microservices. The initiating microservice is the caller, and the receiving one is the callee..
We designed and implemented a robust data collection and analysis pipeline, leveraging our previous work [19, 20, 4]. The pipeline was designed to be data-agnostic and could persistently store data in IBM COS for near-real-time analysis. A Pub/Sub [21] mechanism managed the high data volume, with each microservice publishing telemetry to a Redis Pub/Sub pipeline. Microservices running as containers, managed by Kubernetes, extracted, transformed, and loaded (ETL) the necessary data into persistent storage system.
A sub-pipe connected to the Redis Pub/Sub captured traces in the Zipkin format [22] (with support for extending to formats such as OpenTrace [23] and OpenTelemetry [24]). Telemetry was collected at 1-minute intervals to provide near-real-time information. Due to the large data volume, we filtered the data during collection based on recommendations from the IBM Console team.
The sub-pipe includes an ETL module, extracting essential fields such as timestamps, trace IDs, microservice names, HTTP status codes, and response times from the raw telemetry data. The extracted data are then transformed into arrays capturing computed statistics for each microservice across all production data centers and for caller-callee microservice pairs. These statistics included metrics such as call volumes, average response times, and the distribution of HTTP response codes (see Table I for the complete list).
TABLE I: Metrics generated from the extracted data.
Extracted Fields Metrics
Count of Distinct Services Call Volumes
Duration Average Response Time
Maximum Response Time
Minimum Response Time
Standard Deviation of Response Time
HTTP Status Code Percentage and value of HTTP response codes
Finally, the transformed data were stored as a nested JSON object with indexed timestamps, facilitating efficient time series analysis. The structured data were then used to generate heatmaps providing an intuitive visualization of system health metrics.
II-C Data Processing
Once collected, the raw data needed processing and aggregation into defined intervals suitable for visualization. The data processing phase involved several steps:
Data Aggregation
We combined each 1-minute interval of data stored in IBM Cloud Object Storage (COS) [25] over specific periods (e.g., 24 hours) using IBM Cloud Functions [^5] Replaced by IBM Cloud Code Engine [26] at the time of writing. [27]. A custom action aggregated new files into a combined file with a maximum size of 1 MB per file. After aggregation, functions such as sum, average, maximum, minimum, and standard deviation generated summary statistics for each microservice.
Data Transformation
Aggregated data were transformed into a format suitable for heatmap visualization, converting the aggregated metrics into two-dimensional arrays where rows represented data centers or caller microservices, and columns represented monitored microservices. The transformation included resampling data to desired time intervals (e.g., 15 minutes, hourly) using aggregation functions (e.g., compound mean and standard deviation, see [17] for details).
Data Storage
The processed data were stored in IBM COS, which provided a scalable and secure environment for managing the large datasets the IBM Cloud Console generated. The data retrieval module, written in Python, is connected to the IBM COS API to read objects from the storage bucket within a user-defined time window. The retrieved JSON data was converted into a two-dimensional DataFrame for further processing.
II-D Visualization Design
The core part of the tool involved designing and implementing the heatmap visualization tool using the Plotly Dash library [28, 29].
Figure 2: A heatmap comparing data centers and microservices, useful for assessing whether an issue has a greater impact on data centers.
Figure 3: A heatmap showing caller-callee microservices, useful for identifying issues in the interactions between microservices.
CloudHeatMap visualizes call volumes, response times, and HTTP response codes across the IBM Cloud Console’s microservices and data centers. Figs. 2 and 3 show sample heat maps using actual anonymised data (included with the demo code). Fig. 2 illustrates the relationship between data centers and services, with microservice names on the $x$-axis and data centers on the $y$-axis. Fig. 3 depicts the caller-callee interactions, with callee microservices on the $x$-axis and caller microservices on the $y$-axis. Hovering with a mouse over a specific tile in the heatmap displays a window showing the numeric value of the tile along with its corresponding $x$ and $y$ coordinates, enhancing clarity. Key features include:
Interactive User Interface
The tool offers a highly interactive UI with multiple filters, allowing customization based on graph type (data centers-services or caller-callee pairs), metrics (e.g., call volumes or average response time), and HTTP status codes. Users can switch between viewing absolute values or percentages and filter by value type and range.
Temporal Analysis
Users can play an animation displaying evolution of metrics over time, crucial for identifying trends and episodic issues that static views might miss. The first frame of the animation presents an aggregated view of the selected time period, offering a comprehensive snapshot, with subsequent frames showing behavioral changes over time.
CloudHeatMap  is designed to highlight “hot spots” where likely, indicated by areas of higher colour intensity on the heatmap (e.g., indicating higher call volumes, longer response times, or higher error rates). This helps operators quickly identify and address potential issues.
III Evaluation
We assessed CloudHeatMap through a two-month case study on IBM Cloud Console (discussed in Section II-A) in collaboration with the IBM Cloud Ops team, focusing on its ability to monitor and improve IBM Cloud Console’s performance. For details of the evaluation of use cases, see [17].
III-A Rate Limiting Detection
Rate limiting is a concern in Cloud environments, where microservices must operate within specified limits to prevent service degradation. CloudHeatMap analyzed HTTP 429 “Too Many Requests” response codes across microservices revealing that one microservice consistently exceeded acceptable rate limits across data centers, with peak values reaching 40% in some time intervals. This visualization allowed the Ops team to identify and address the issue promptly.
III-B Detecting Components’ Errors
Server-side errors (HTTP 5XX status codes) critically impact user experience. CloudHeatMap  enabled the Ops team to efficiently detect and address these errors. For example, it revealed services with a 100% error rate, leading to immediate corrective actions, including removing obsolete services and rectifying active ones. The team prioritized services with significant call rates and planned to address lower-impact ones subsequently.
Not all endpoints return standard HTTP status codes. Some services use non-standard codes, like -1, for custom protocol interactions. CloudHeatMap successfully visualized both HTTP and non-HTTP traffic, providing a comprehensive system view.
III-C Detecting Performance Degradation
User satisfaction is closely tied to interface responsiveness, with response times over 2 seconds degrading experience [30]. CloudHeatMap  identified (using caller-callee view) microservices contributing to latency issues, such as “Dashboard-broker” and “Preferences” interacting with a “Cloudant” database instance. Early detection allowed the Ops team to intervene before issues escalated into a critical failure.
III-D Re-architecting Hot-spots
Analysis of the “Datalayer” microservice using CloudHeatMap  identified inefficient calls, leading to improved caching strategies. The team re-architected the system, replacing the caching mechanism with an enhanced “GraphQL” layer, reducing call volumes and increasing flexibility for the front-end team.
III-E Cost Savings
CloudHeatMap contributed to cost-saving initiatives by identifying infrequently used microservices with low traffic, deployable on smaller, more cost-effective Kubernetes clusters, allowing the IBM Ops team to explore operational cost reductions through optimized resource deployment.
IV Discussion
CloudHeatMap  equipped the IBM Ops team with actionable insights that traditional monitoring tools could not, revealing hidden issues and distinguishing between persistent and episodic problems. This proactive monitoring approach provides a more comprehensive view of system health, unlike the retrospective analyses of other platforms like Datadog and Dynatrace. Additionally, CloudHeatMap bridges the gaps left by machine-learning-based tools, which, despite their sophistication, often struggle with exploratory analysis [31].
These use cases align with some of Musa’s Operational Profiles analysis [32], where observing client usage helps identify problematic components for testing and re-architecting, especially those frequently accessed or with high response times.
CloudHeatMap answered RQ1 by providing visualizations that allowed operators to monitor three of the four golden signals recommended by Google SRE — traffic, errors, and latency. These visualizations enabled the Cloud Ops team to quickly identify and address performance issues, proving the visualizations to be actionable and insightful. For RQ2, the tool analyzed call volumes and workload patterns across different data centers, identifying bottlenecks and components struggling under varying workloads. This ensured operators could better understand components behaviours under dynamic conditions. For RQ3, CloudHeatMap  allowed near-real-time issue detection by enabling operators to monitor critical metrics as they evolved. The heatmap facilitated early identification of performance degradation, error spikes, and anomalies, allowing proactive issue resolution. For RQ4, the tool’s insights directly impacted the maintenance process. The visualized data guided decisions like resource scaling and re-architecting high-traffic or error-prone components, improving system reliability and efficiency. These insights were validated through real-world use cases with the IBM Ops team.
V Threats to Validity
The primary validity concerns, classified as per [33, 34], are as follows.
Construct Validity
The tool was iteratively developed with feedback from the IBM Ops team and tested in production for two months, ensuring alignment with evaluation goals and mitigating potential threats to construct validity.
Internal Validity
Rigorous testing and verification of the data were employed to minimize experimental errors.
External Validity
The monitoring tool is not limited to one software system. While not tested on other products, it is designed to be generic and applicable to any LCS.
VI Conclusion
We proposed CloudHeatMap, a heatmap-based tool to monitor LCS health, using the IBM Cloud Console as a software under study. The tool represents a significant advance in Cloud monitoring by providing actionable insights in real-time, allowing operators to make informed decisions about scaling, re-architecting, and cost-saving measures. Its unique approach to visualizing both HTTP and non-HTTP interactions makes it adaptable to various Cloud environments. We believe CloudHeatMap  will interest practitioners and academics and contribute to monitoring complex large-scale software systems.
Moving forward, we will focus on deriving the fourth Golden Rule of Monitoring — saturation — to provide a more complete picture of system health. Aggregating caller-callee pairs into call chains for graph-based root cause analysis will further enhance the tool’s diagnostic capabilities.